OpenStudy (zab505):

According to the rational root theorem, which of the following are possible roots of the polynomial function below? F(x) = 8x3 - 3x2 + 5x + 15 A. 5 B. 4 C. -1/4 D. -3 E. 5/2 F. 2/3

OpenStudy (amistre64):

how do we use the rational roots thrm?

OpenStudy (zab505):

Not sure?

OpenStudy (amistre64):

then you may want to read what the rational roots thrm is, give me a definition of it and lets see what part of it your not grasping.

OpenStudy (zab505):

The number by which it's all divided?

OpenStudy (amistre64):

hmm, not quite. my recollection of the rrt is that the only possible rational roots that a poly can have can be determined from a pool of options created from the factors of the constant term divided by the factors of coefficient of the highest degree in the poly.

OpenStudy (amistre64):

the factors of your constant term: 15, are 1,3,5,15 the factors of highest degree term: 8, are 1,2,4,8 so the rrt says that if we are to have any rational root, it must be of the form:\[\pm\frac{1,3,5,15}{1,2,4,8}\]

OpenStudy (amistre64):

now the question is, what can we form from these? A. 5/1 B. 4/1 C. -1/4 D. -3/1 E. 5/2 F. 2/3

OpenStudy (zab505):

What do you mean by form?

OpenStudy (amistre64):

pick a number from the top, and divide it by a number from the bottom. those are the only numbers we can use to form a fraction with

OpenStudy (amistre64):

for instance:\[\frac{4}{7}\not \in\frac{1,3,5,15}{1,2,4,8}\] since we cannot form 4/7 out of the given options

OpenStudy (amistre64):

but 5/1 is an option since \[\frac{5}{1} \in\frac{1,3,\color{red}5,15}{\color{red}1,2,4,8}\]

OpenStudy (zab505):

That would be the only one possible?

OpenStudy (amistre64):

i see 2 that are not possible, and the rest are

OpenStudy (amistre64):

see if you can make a stated option from the given pool of numbers

OpenStudy (anonymous):

please help @amistre64

OpenStudy (zab505):

Would it be everything except B and F?

OpenStudy (zab505):

@amistre64